# B2MAPO: A Batch-by-Batch Multi-Agent Policy Optimization to Balance Performance and Efficiency

### v1

@@ 3.2.1 Batch Sequence Generation Layer
3.2.2 Batch by Batch Multi-Agent Policy Optimization Layer
@@
The CTDE-based joint policy $\boldsymbol{\pi}_{ind}$ is trained with different experiences gathered by joint policy $\boldsymbol{\pi}$ that is produced by B2MAPO scheme. Meanwhile, according to theorem 3, $\boldsymbol{\pi}_{ind}$ is derived from $\boldsymbol{\pi}$ by periodically minimize the KL divergence between $\boldsymbol{\pi}_{ind}$ and $\boldsymbol{\pi}$.(Proofs can be found in the Appendix). $\boldsymbol{\pi}_{ind}$ inferences efficiently with lightweight networks and satisfies CTDE principle, $\boldsymbol{\pi}$ supports B2MAPO scheme and guarantees optimal performance.
@@
B2MAPO framework provides a plug-and-play, modulized and flexible architecture, which can conveniently integrate third-party models with few or even without modification to exploit their merits. "CTDE-based joint policy" module in this layer can be implemented by various MARL methods, such as Q-learning-based methods[30, 35, 39], policy-based methods[41], and actor-critic methods[44].
@@ 4 DAG-based B2MAPO Algorithm
DAG-based B2MAPO algorithm is a carefully designed implementation of the universal B2MAPO framework. The upper layer contains a DAG critic and a DAG generator trained by PPO algorithm[34]. The DAG generator constructs DAGs of agents to produce optimal batch sequence. In the lower layer, we design a network of policy for an agent, which can be shared among all agents. Joint policy $\boldsymbol{\pi}$ is trained based B2MAPO scheme with optimal batch sequence, while the derived joint policy $\boldsymbol{\pi}_{ind}$ is updated by MAPPO. Figure 3 presents the network architecture of the DAG-based B2MAPO algorithm.
@@ 4.1 Hierarchical DAG Generation
is employed in DAG generator $\pi_{\theta}^{dag}$ to analyze the interdependence between agents for more accurate spinning subgraph DAG $G$ generation. The upper layer outputs the batch sequence $(B,\prec)$ through topological sorting of DAG $G$. Since the optimal batch sequence $(B^{opt},\prec)$ is closely related to the dependence among agents,
During centralized training, at time step $t$, for each agent $i$, DAG generator encodes the observation $o_{t}^{i}\!\in\!\boldsymbol{o}$ into the historical trajectory $\tau_{t}^{i}\!\in\!\boldsymbol{\tau}_{t}$ by multilayer perceptron(MLP) and recurrent neural network (RNN) module. While the historical trajectory $\tau^{i}_{t}$ of each agent $i$ is encoded, the query vectors $q^{i}=\textbf{W}_{q}\tau^{i}_{t}$ , the key vectors $k^{i}=\textbf{W}_{k}\tau^{i}_{t}$ and the value vectors $v^{i}=\textbf{W}_{v}\tau^{i}_{t}$ are calculated as well.
Through attention mechanism, we derive an estimated graph $G_{T}=\{G_{T}^{i}\}$ which represents interdependence between agents, and $G_{T}^{i}$ signifies the degree of dependency between agent $i$ and other agents.
$G_{T}^{i}=\text{softmax}\left[\frac{q_{i}^{\top}k_{1}}{\sqrt{d_{k}}},\frac{q_{i}^{\top}k_{2}}{\sqrt{d_{k}}},\ldots,\frac{q_{i}^{\top}k_{n}}{\sqrt{d_{k}}}\right]$ (5)
In addition, we set $g_{T}^{ij}=0$, if $||G_{T}^{ij}||<\delta$, where $\delta>0$ denotes a fixed threshold.
Then, the estimated graph $G_{T}$ is encoded through an MLP to obtain an abstract representation $Z_{T}=\{v^{i}{G_{T}^{i}}\}$ of the current interdependence between agents. Subsequently, the abstract representation $Z_{T}$ of dependency relationships is embedded to generate and initialize the parameters of DAG construction module. Our DAG construction module is implemented based on [4] DAG construction module outputs the spinning subgraph DAG $G$ based on the estimated graph $G_{T}$. The batch sequence $(B,\prec)$ for the future $T$ steps is obtained by topological sorting of the spinning subgraph DAG $G$.
DAG Critic is employed to estimate the advantage function $\hat{A}^{dag}$ with $(B,\prec)$ from time step $t$ to $t+T$; Similar to DAG Generator, at time step $t$, DAG Critic encodes joint observation $\boldsymbol{o}_{t}$ into joint historical trajectory $\boldsymbol{\tau}_{t}$, through MLP and RNN module. Then, DAG Critic generates the estimated state value $V_{\pi}^{dag}(s_{t})$ based on the embedding of the global state $s_{t}$ and joint historical trajectory $\boldsymbol{\tau}_{t}$. Finally, DAG Critic, combined with Generalized Advantage Estimation (GAE)[33] module for more precise estimation, approximates the advantage function $\hat{A}^{dag}$ for the future $T$ steps:
@@ 5.1 Performance Analysis
Google Football Research(GRF) In this section, we test our B2MAPO on two challenging GRF scenarios: academy_counterattack_easy and academy_counterattack_hard scenarios, which pose challenges for agents in discovering complex coordination behaviors. We compare performance against baselines in Figure 5. MAPPO struggles to learn sophisticated policies in all scenarios. Due to heuristic order sequence generation, as the difficulty of the experiment increases, the performance of the A2PO algorithm deteriorates, and the instability increases, which is reflected in the high variance and low winning rate in the academy_counterattack_hard scenario. Our B2MAPO outperforms all baselines in all the scenarios with a stable training process, demonstrating superior performance in complex cooperative tasks. Especially in the difficult academy_counterattack_hard scenario, B2MAPO has a higher and more stable performance, far better than A2PO, mainly reflected in small variance and high winning rate.
@@ 7 Conclusion
In conclusion, we introduce the B2MAPO scheme with the off-policy correction method, which theoretically guarantees the monotonic incrementally tightened bound. Then, a universal modulized plug-and-play B2MAPO hierarchical framework, which satisfies CTDE principle, is designed to conveniently integrate any MARL models to fully exploit and merge their merits. Based on the hierarchical B2MAPO framework, a deliberative implementation, called DAG-based B2MAPO algorithm is proposed. The upper layer employs PPO algorithm with attention mechanism to generate DAGs of agents which are used to produce optimal batch sequence through topological sorting. Based on the theoretical proof, the lower layer trains two joint policies in parallel and minimizes KL divergence between them periodically. The final derived joint policy achieves high efficiency while ensuring algorithm optimality. Experiments consistently showcase superior performance compared to baseline methods while significantly improving the efficiency during the training and execution phase. Overall, B2MAPO presents a promising direction in multi-agent decision-making, offering a principled approach that balances performance, efficiency, and adaptability.

### v2

@@ 3.2.1 Batch Sequence Generation Layer
3.2.2 Batch by Batch Policy Optimization Layer
@@
$\boldsymbol{\pi}_{ind}$, which is composed of independent policies of all agents, inferences efficiently with lightweight networks and satisfies CTDE principle. $\boldsymbol{\pi}$ supports B2MAPO scheme and guarantees optimal performance. The CTDE-based derived joint policy $\boldsymbol{\pi}_{ind}$ is trained with different experiences gathered by joint policy $\boldsymbol{\pi}$ that is produced by B2MAPO scheme. Meanwhile, according to theorem 3, $\boldsymbol{\pi}_{ind}$ is derived from $\boldsymbol{\pi}$ by periodically minimize the KL divergence between $\boldsymbol{\pi}_{ind}$ and $\boldsymbol{\pi}$. (Proofs can be found in [zhang2024b2mapo])
@@
B2MAPO framework provides a universal modulized plug-and-play architecture, which can conveniently integrate third-party models with few or even without modification to exploit their merits. "CTDE-based joint policy" module in this layer can be implemented by various MARL methods, such as Q-learning-based methods[30, 35, 39], policy-based methods[41], and actor-critic methods[44].
@@ 4 DAG-based B2MAPO Algorithm
DAG-based B2MAPO algorithm is a carefully designed implementation of the universal B2MAPO framework. The upper layer contains a DAG critic and a DAG generator trained by PPO algorithm[34]. The DAG generator constructs DAGs of agents to produce optimal batch sequence. In the lower layer, we design a policy network for an agent, which can be shared among all agents. Joint policy $\boldsymbol{\pi}$ is trained based B2MAPO scheme with optimal batch sequence, while the derived joint policy $\boldsymbol{\pi}_{ind}$ is updated by MAPPO. Figure 3 presents the network architecture of the DAG-based B2MAPO algorithm.
@@ 4.1 Hierarchical DAG Generation
The DAG generator utilizes attention mechanism build agent interdependent graph by encoding historical trajectories every $T$ time steps. DAG construction module periodically collects the graph to produce its spinning subgraph (DAG of agents). The optimal batch sequence is generated based on the DAG by topological sorting.
During centralized training, at time step $t$, for each agent $i$, DAG generator encodes the observation $o_{t}^{i}\!\in\!\boldsymbol{o}$ into the historical trajectory $\tau_{t}^{i}\!\in\!\boldsymbol{\tau}_{t}$ by multilayer perceptron(MLP) and recurrent neural network (RNN) module. While the historical trajectory $\tau^{i}_{t}$ of each agent $i$ is encoded, the query vectors $q^{i}=\textbf{W}_{q}\tau^{i}_{t}$ , the key vectors $k^{i}=\textbf{W}_{k}\tau^{i}_{t}$ and the value vectors $v^{i}=\textbf{W}_{v}\tau^{i}_{t}$ are calculated as well. Then, the graph $G_{T}=[G^{1}_{T},G^{2}_{T},\ldots,G^{n}_{T}]^{\top}$ represented the interdependence between agents is derived, where $G^{i}_{T}$ signifies the degree to which agent $i$ depend on other agents.
$G^{i}_{T}=[g_{T}^{ij}]=\text{softmax}\left[\frac{q_{i}^{\top}k_{1}}{\sqrt{d_{k}}},\frac{q_{i}^{\top}k_{2}}{\sqrt{d_{k}}},\ldots,\frac{q_{i}^{\top}k_{n}}{\sqrt{d_{k}}}\right]$ (5)
where $g_{T}^{ii}=0$. In addition, if $g_{T}^{ij}<\delta$, $g_{T}^{ij}=0$, where $\delta>0$ is a hyper-parameter indicated the minimum degree of interdependence.
The graph $G_{T}$ is embedded by an MLP to obtain an abstract representation $Z_{T}=MLP(v^{i}{G^{i}_{T}})$, which is utilized to generate and initialize the parameters of DAG construction module. The DAG construction module is implemented based on [4] DAG construction module outputs the spinning subgraph $G$ based on $G_{T}$. The batch sequence $(B,\prec)$ for the future $T$ steps is obtained by topological sorting of the spinning subgraph DAG $G$.
DAG Critic is employed to estimate the advantage function $\hat{A}^{dag}$ with $(B,\prec)$ from time step $t$ to $t+T$; Similar to DAG Generator, at time step $t$, DAG Critic encodes joint observation $\boldsymbol{o}_{t}$ into joint historical trajectory $\boldsymbol{\tau}_{t}$, through MLP and RNN module. DAG Critic generates the estimated state value $V_{\phi}^{dag}(s_{t})$ based on the embedding of the global state $s_{t}$ and joint historical trajectory $\boldsymbol{\tau}_{t}$. Finally, DAG Critic, combined with Generalized Advantage Estimation (GAE)[33] module for more precise estimation, approximates the advantage function $\hat{A}^{dag}$ for the future $T$ steps:
@@ 5.1 Performance Analysis
Google Football Research(GRF) In this section, we test B2MAPO on two challenging GRF scenarios: academy_counterattack_easy and academy_counterattack_hard scenarios, which pose challenges for agents in discovering complex coordination behaviors. We compare performance against baselines in Figure 5. MAPPO struggles to learn sophisticated policies in all scenarios. Due to heuristic order sequence generation, as the difficulty of the experiment increases, the performance of the A2PO algorithm deteriorates, and the instability increases, which is reflected in the high variance and low winning rate in the academy_counterattack_hard scenario. Our B2MAPO outperforms all baselines in all scenarios with a stable training process, demonstrating superior performance in complex cooperative tasks. Especially in the difficult academy_counterattack_hard scenario, B2MAPO has a higher and more stable performance, far better than A2PO, mainly reflected in small variance and high winning rate.
@@ 7 Conclusion
In conclusion, we introduce the B2MAPO scheme with the off-policy correction method, which theoretically guarantees the monotonic incrementally tightened bound. Then, a universal modulized plug-and-play B2MAPO hierarchical framework, which satisfies CTDE principle, is designed to conveniently integrate any MARL models to fully exploit and merge their merits. Based on B2MAPO framework, a deliberative implementation, called DAG-based B2MAPO algorithm is proposed. The upper layer employs PPO algorithm with attention mechanism to generate DAGs of agents and produce optimal batch sequence through topological sorting. Based on the theorem, the lower layer trains two joint policies in parallel and minimizes KL divergence between them periodically. The final derived joint policy achieves high efficiency while ensuring algorithm optimality. Experiments consistently show superior performance compared to baseline methods while significantly improving the efficiency during the training and execution phase. Overall, B2MAPO presents a promising direction in multi-agent decision-making, offering a principled approach that balances performance, efficiency, and adaptability.